Economics

Assume Ali and Barb specialize in their comparative advantage good, and then agree to trade 3 loafs of bread for 12 jars of jam. How many loafs of bread and jars of jam do they each have now?
Answer #1

For Ali,


Opportunity cost of producing loafs of bread = 20/2 = 2 jars of Jam


For Barb,


Opportunity cost of producing loafs of bread = 36/6= 6 jars of Jam


Since Ali has lower opportunity cost of producing bread than that of Barb, so Ali has comparative advantage in producing bread and Barb has comparative advantage in producing Jam.


Now after specialization, 3 loafs o bread are traded with 12 jars of jam,


Then;


Ali will have:


7 loafs of bread and 12 jars of jam


Barb will have:


3 loafs of bread and 24 jars of jam


answered by: hasheco